Keniyahwethepeople:

Find the volume of a cone with a base diameter of 12 in and a height of 11 in .

Keniyahwethepeople:

Need help on.

darkknight:

Do you know the formula for a volume of a cone?

darkknight:

1/3)(πhr^2 is the formula

darkknight:

Where r is radius. and h is height

darkknight:

Now all you need to do is to plug and chug. radius = 1/2(diameter) = 6 and height = 11

Keniyahwethepeople:

V= pi r^2 h/3

darkknight:

Yes, that is the correct formula

Keniyahwethepeople:

ok so what next

darkknight:

So next you have to plug in the values. radius = 1/2(diameter) = 6 and height = 11

Keniyahwethepeople:

oh okay

darkknight:

So what did you get?

Keniyahwethepeople:

414.69 right

darkknight:

Yes

darkknight:

Good job Keep it up :)

Keniyahwethepeople:

Thx for the help!

darkknight:

Yeah, np

darkknight:

click the best response button and fan if ya think I have helped. Hope you have a nice day/night.